OpenStudy (anonymous):

Verify the identity. (8 points) cot^2 x/csc x + 1 = 1 - sinx/ sin x i really need help

OpenStudy (mathmath333):

is it this \(\large \begin{align} \color{black}{ \dfrac{\cot^2 x}{\csc x} + 1 = 1 - \dfrac{\sin x}{ \sin x} \hspace{.33em}\\~\\ or ~~this \hspace{.33em}\\~\\ \dfrac{\cot^2 x}{\csc x} + 1 = \dfrac{1-\sin x}{ \sin x} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

let me attach a photo

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{\cot^2x}{\csc x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x}{\csc x-1}\times \dfrac{\csc x+1}{\csc x+1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\csc^2 x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\csc^2 x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\cot^2x} \hspace{.33em}\\~\\ =\csc x+1 \hspace{.33em}\\~\\ =\dfrac{1}{\sin x}+1 \hspace{.33em}\\~\\ }\end{align}\) see if it makes sense

OpenStudy (anonymous):

thats only part of it right?

OpenStudy (anonymous):

csc x+1 denominator

OpenStudy (anonymous):

can u show me step by step to get the answer and then what the answer would be

OpenStudy (mathmath333):

yes 99% the last part can be written as \(\large \begin{align} \color{black}{ =\dfrac{1}{\sin x}+1 \hspace{.33em}\\~\\ =\dfrac{1+\sin x}{\sin x} \hspace{.33em}\\~\\ }\end{align}\) now see if it does match with th question given

OpenStudy (anonymous):

well what I'm trying to understand is why you switched csc x + 1 to a negative 1

OpenStudy (mathmath333):

u mean this red part \(\large \begin{align} \color{black}{ \dfrac{\cot^2x}{\csc x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x}{\csc x-1}\times \color{red }{\dfrac{\csc x+1}{\csc x+1} }\hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\csc^2 x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\csc^2 x-1} \hspace{.33em}\\~\\ =\dfrac{\cot^2x\times (\csc x+1)}{\cot^2x} \hspace{.33em}\\~\\ =\csc x+1 \hspace{.33em}\\~\\ =\dfrac{1}{\sin x}+1 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

yes and what happened to the right of the problem????

OpenStudy (anonymous):

1 - sinx and sin x

OpenStudy (mathmath333):

no thing happened the question says verify the identity as the last part doesnt matches the one given in the question, the given identity is \(\bf false\) amd about the red part do u agree that \(\large \dfrac{\csc x+1}{\csc x+1}=1\) ?

OpenStudy (anonymous):

yes so what should i submit as an an answer and they want the work

OpenStudy (mathmath333):

if they want work , show as i did upto \(\dfrac{1+\sin x}{\sin x}\) and write that \(LHS\neq RHS\) so its false

OpenStudy (mathmath333):

that red part i did was to rationalise the denominator

OpenStudy (anonymous):

thank you for everything